You want to calculate the cost of debt for a firm with bank debt and long-term bond outstanding. The $100 million revolving line of credit is priced at 1ibor + 250 basis points and with $25 million utilized. The company has a $150 million 5-year bond that pays interest semi-annually. The bond has an 8% coupon$1,000 par value and is currently priced at 98.5 as a percent of par. The Libor reference rate is currently 0.50%. What is the blended average cost of debt for the firm.

a. 6.22%

b. 4.02%

c. 7.61%

d. 3.71%
Answer #1

blended average cost of debt = (weight of bank loan * rate on bank loan) + (weight of bonds * yield on bonds)

value of bank loan = loan utilized = $25,000,000

value of bonds = face value * price as a percent of par = $150,000,000 * 98.5% = $147,750,000

total value of debt = $25,000,000 + $147,750,000 = $172,750,000

weight of bank loan =  $25,000,000 / $172,750,000 = 0.145

weight of bonds =  $147,750,000 / $172,750,000 = 0.855

rate on bank loan = LIBOR + 2.5% = 0.5% + 2.5% = 3%

yield on bonds is calculated using RATE function in Excel with these inputs :

nper = 5*2 (10 years to maturity with 2 semiannual coupon payments each year)

pmt = 1000 * 8% / 2 (semiannual coupon payment = face value * annual coupon rate / 2. This is a positive figure as it is an inflow to the bondholder)

pv = -985 (current bond price = face value * price as % of par / 100 = $1000 * 98.5 / 100 = $985. This is a negative figure as it is an outflow to the buyer of the bond)

fv = 1000 (face value of the bond receivable on maturity. This is a positive figure as it is an inflow to the bondholder)

The RATE is calculated to be 4.19%. This is the semiannual YTM. To calculate the annual YTM, we multiply by 2. Annual YTM is 8.37%

A1 X f =RATE(5*2,1000*8%/2,-985, 1000)*2 B C D F G A 8.37%! |

blended average cost of debt = (weight of bank loan * rate on bank loan) + (weight of bonds * yield on bonds)

blended average cost of debt = (0.145 * 3%) + (0.855 * 8.37%)

blended average cost of debt = 7.61%